# Planet formation theory: an overview

### v1

@@ Abstract
The standard model for planet formation is a bottom-up process in which the origin of rocky and gaseous planets can be traced back to the collision of micron-sized dust grains within the gas-rich environment of protoplanetary disks. Key milestones along the way include disk formation, grain growth, planetesimal formation, core growth, gas accretion, and planetary system evolution. I provide an introductory overview of planet formation, emphasizing the main ideas and reviewing current theoretical understanding. Many of the phases of planet formation have a well-developed physical understanding, though the complexity of the problem means that few can be quantitatively modeled with complete confidence. Transformative advances in disk imaging and exoplanet detection provide the first direct information on the initial conditions for planet formation, and have motivated new formation models that are faster, more efficient, and lead to a more diverse set of architectures than their Solar System inspired forebears. Much remains to be learned, and I close with a personal, incomplete list, of open problems.
@@ 1 Introduction
This review provides an elementary introduction to this standard model, and serves as an overview of the following Chapters which discuss planet formation theory in detail. I introduce the physical processes relevant for planet formation, and summarize how they fit together in our consensus understanding of how planets form. Although a plausible outline of planet formation theory has been known for decades, existing models are surely not the last word: despite the genuinely transformative advances in exoplanet detection and disk imaging there remain few direct observations of planet formation processes. It is conceivable that are fundamental gaps in our current theory. Even within the established paradigm, some phases are known to be poorly understood, and I conclude by listing some open questions that I hope to see resolved soon.
@@ 2 Disk formation and evolution
Although limited grain growth may occur during star formation prior to disk formation (Kwon et al. 2009; Marchand et al. 2023), the vast majority of planet formation processes occur within protoplanetary disks. Disks are flattened near-equilibrium fluid structures, whose basic physical properties are determined observationally to varying (mostly low) levels of precision. Typical stellar accretion rates scale roughly as $\dot{M}\sim 10^{-8}(M_{*}/M_{\odot})^{2}M_{\odot}{\rm yr}^{-1}$, with a measurement precision of $\sim 0.35\ {\rm dex}$ and a spread at fixed stellar mass $M_{*}$ of more than an order of magnitude (Manara et al. 2023). The disk lifetime is around 3 Myr (Haisch et al. 2001), though the absolute ages of Young Stellar Objects are poorly determined (Soderblom et al. 2014). The relative ages of young clusters are more reliable. Disk masses can be estimated starting from various observables, including the dust continuum flux (Ansdell et al. 2016; Tychoniec et al. 2020), molecular line emission in commonly observed sub-mm tracers (CO and its isotopologues, sometimes corrected using additional species such as N2H+; Trapman et al. 2022), HD emission (McClure et al. 2016), and (at the high mass end) rotation curves (Martire et al. 2024; Andrews et al. 2024). All of these methods have drawbacks – either from limited applicability or from possibly large uncertainties – such that a consensus empirical model for disk evolution remains elusive. Planet formation theorists often assume as initial conditions a disk gas surface density profile that scales with radius as $\Sigma\propto r^{-1}$, an initial gas to dust ratio of $10^{2}$, and a mass within 100 au of $10^{-3}-0.1\ M_{*}$, but this is more a matter of art and convention than anything based on solid data. Despite its well-known shortcomings, the Minimum Mass Solar Nebula (Weidenschilling 1977b; Hayashi 1981) also lives on in some computers.
@@
where $\rho$ is the density, $\Omega_{\rm K}$ the Keplerian angular velocity, and we have assumed that $z\ll r$, the radial distance to the star. For a vertically isothermal disk with sound speed $c_{s}$, $P=\rho c_{s}^{2}$ and the gas density declines as a gaussian,
@@
Except in cases of unusually high accretion rates, or when close to the star, the heating of protoplanetary disks occurs primarily from the interception and re-radiation of stellar luminosity (Kenyon and Hartmann 1987). Irradiation leads to a radial temperature profile approximately given by a power-law, $T\propto r^{-3/7}$ (Chiang and Goldreich 1997), and a disk that flares, such that $h/r$ increases with radius. Typically, $0.10.03<h/r<$ 0.1. In the radial direction, the force balance involves gravity, centrifugal force, and a pressure gradient,
@@
The pressure gradient term is order $(h/r)^{2}$ smaller than the gravitational one, so to leading order $\Omega(r)=\Omega_{\rm K}$. The small correction – which usually but not always leads to sub-Keplerian rotation – is nonetheless extremely important for planet formation. The dynamics of dust particles (Whipple 1972; Weidenschilling 1977a), the formation of planetesimals (Youdin and Goodman 2005), and the interpretation of observed disk sub-structures (Andrews 2020) are all thought to hinge on this $\sim$1deviation from strict Keplerian motion.
Figure 1: An illustration of the main processes that lead to evolution of the gaseous component of protoplanetary disks.
@@
  • Processes that redistribute angular momentum within the disk, while conserving disk mass (up to the fraction of gas that is physically accreted by the star). Lesur et al. (2023) review the long laundry list of possible processes, most of which are associated with linear instabilities of disk flow. Self-gravity (Toomre 1964; Kratter and Lodato 2016), the magnetorotational instability (Balbus and Hawley 1991) as (strongly) modified by non-ideal magnetohydrodynamic effects, and the Vertical Shear Instability (Nelson et al. 2013) are the leading suspects.
  • Loss of mass and / or angular momentum through disk winds. Photoevaporation (Hollenbach et al. 1994) – in which the surface of the disk is heated by high-energy photons to the point where it escapes in a thermal wind – is guaranteed to occur at some level (Alexander et al. 2014; Ercolano and Pascucci 2017). It is a pure mass-loss process. Magnetized winds (Blandford and Payne 1982), on the other hand, remove gas that has more angular momentum than the disk at the launch point, simultaneously dispersing the disk while driving accretion (Bai and Stone 2013a, b)
Figure 1 illustrates these basic possibilities. In addition to evolution, many of these processes can also plausibly generate rings, spirals, and other sub-structures in disks.
The state-of-the-art in direct numerical simulation (Xu and Kunz 2021; Mauxion et al. 2024) allows disks to be followed for about $10^{5}\ {\rm yr}$, roughly the duration of the Class 0 phase of Young Stellar Object evolution. Longer duration evolution, and studies that seek to compare theory against observed populations of disks, necessarily require simpler models. The minimal goal of one-dimensional models is to predict $\Sigma(r,t)$, though more complete treatments may also evolve the disk central temperature (this is essential for modeling disk outbursts; Bell and Lin 1994; Armitage et al. 2001; Zhu et al. 2009). The vertical magnetic flux threading the disk, $B_{z}(r,t)$, is another important dynamical quantity that in principle should be evolved in models that invoke magnetized winds or the magnetorotational instability. The simplest case is a disk that evolves solely due to local turbulent redistribution of angular momentum. The appropriate one-dimensional description is a diffusion equation (Lynden-Bell and Pringle 1974; Pringle 1981),
@@
in which $\nu$ is an effective viscosity that is generated by turbulence. If $\nu$ is approximated by a time-independent power-law function of radius, this equation admits a similarity solution (a time-varying exponentially truncated power-law for the surface density) that has been widely used for disk modeling (e.g. Hartmann et al. 1998). More generally, the viscosity can be written in terms of the local scale height $h$ and disk sound speed $c_{s}$ as (Shakura and Sunyaev 1973),
@@
with $\alpha$ being a dimensionless parameter that characterizes the efficiency of angular momentum transport. Wind-driven evolution can be modeled within an analogous framework, by adding sink and advective terms to the right hand side of equation (5) that represent the effects of mass and angular momentum loss on the underlying disk. The detailed form of the additional terms depends on the choice of wind model; Tabone et al. (2022) derive a generalization of the Lynden-Bell and Pringle (1974) solution for one choice that may be generic enough to be useful for population modeling. Some effects due to the infall of gas onto the disk are also amenable to a (numerical) one-dimensional description (Lin and Pringle 1990).
@@
Determining the relative importance of these evolution processes, from either theoretical or observational arguments, is a major challenge. It has often been assumed that the ordering from top to bottom in Figure 1 maps onto a temporal sequence, in which disks form from infall on a time scale of $\sim 10^{5}\ {\rm yr}$, evolve initially as massive self-gravitating structures and then more slowly under the action of viscosity and winds, and are finally dispersed by photoevaporation. This remains a plausible paradigm. Observations of molecular line broadening (Flaherty et al. 2017), and of the thickness of dust layers (Villenave et al. 2022), imply low levels of turbulence in many (though not all; Flaherty et al. 2024) of the best-studied disks. The inferred values of $\alpha$ in the “weakly turbulent” disks – typically $\alpha\sim 10^{-4}$ or lower – are broadly consistent with strongly damped MHD turbulence (described as “dead zone” models; Gammie 1996) and weak purely hydrodynamic driving mechanisms. As a consequence, magnetized disk winds rather than turbulence are assumed to be dominant in many recent models. There remains room, however, for larger deviations from conventional wisdom, such as models in which infall is dynamically important at late as well as early times.
@@
A primary result of ALMA (Atacama Large Millimeter/submillimeter Array) has been the discovery that protoplanetary disks typically show pronounced substructure, in the form of departures from the simplest model of an axisymmetric disk with a monotonically decreasing profile of intensity (Andrews 2020). Substructure can be detected most economically in dust continuum imaging, but is also seen in molecular line data that directly traces the gas. Rings, crescent-shaped features, and spirals are all observed, roughly in decreasing order of frequency. Substructure is also detected in scattered light images of disks (Avenhaus et al. 2018), most prominently around more massive stars. Scattered light substructure arises from small grains in the disk atmosphere, and is often harder to interpret than sub-mm images that trace regions closer to the mid-plane. Necessarily, all of the observed substructure involves orbital radii and spatial scales $>{\rm au}$. There are robust theoretical reasons for the existence of cavities and rings at sub-au scales, including from the interaction of a stellar magnetosphere with the disk (Zhu et al. 2024) and from the onset of stronger turbulence due to the magnetorotational instability at temperatures where the alkali metals are thermally ionized (Gammie 1996). These unobserved structures could be central to our understanding of close-in planet formation (Chatterjee and Tan 2014).
@@
The proximate cause of currently observed substructure in dust continuum images is understood to be the aerodynamic concentration of particles in local gas pressure maxima (or, in weaker cases, local pressure enhancements over a smooth background). This process – often described imprecisely as “trapping” – was foreseen by Whipple (1972) (if you haven’t already, look at his amazingly prescient Figure 1). The pressure maxima can be in form of axisymmetric rings (Haghighipour and Boss 2003), as in the original Whipple proposal, vortices (Barge and Sommeria 1995), or spiral waves (Rice et al. 2004), matching the main classes of observed structures. Ring-like dust structures are also predicted to form without the need for underlying gas perturbations in the vicinity of condensation / sublimation fronts (Zhang et al. 2015), but this effect does not appear to be responsible for most of the substructure that is currently being seen.
@@
While the immediate cause of dust substructure is reasonably clear, the same cannot be said for why gas disks develop pressure maxima in the first place. A popular hypothesis is that the source of substructure is gravitational perturbations from already-formed planets. Planets launch spiral density waves, and open gaps whose edges – which are approximately axisymmetric pressure maxima – can become unstable to vortex generation through the Rossby Wave Instability (Lovelace et al. 1999). Planet-disk interaction models generate disk substructure that is often a good match to what is actually observed (Zhang et al. 2018), and are surely responsible for some substructure. That said, simulations of disk turbulence and magnetized winds also commonly lead to the formation of rings (“zonal flows”; Johansen et al. 2009) and vortices, while self-gravity of course forms spirals. As a result, the interpretation of substructure is uncertain. For the purposes of this review, the key takeaway is that the observations of substructure have profound implications for planet formation irrespective of their interpretation. If ringed disks almost always host planets, that means that a substantial population of ice or gas giant planets is able to form surprisingly quickly at orbital separations that are as large as $\sim 50~{}{\rm au}$. If substructure is instead a byproduct of disk transport processes, then the observed concentration of solids that results modifies the evolution of all subsequent phases of planetary growth.

### v2

@@ Abstract
The standard model for planet formation is a bottom-up process in which the origin of rocky and gaseous planets can be traced back to the collision of micron-sized dust grains within the gas-rich environment of protoplanetary disks. Key milestones along the way include disk formation, grain growth, planetesimal formation, core growth, gas accretion, and planetary system evolution. I provide an introductory overview of planet formation, emphasizing the main ideas and reviewing current theoretical understanding. Many of the phases of planet formation have a well-developed physical understanding, though the complexity of the problem means that few can be quantitatively modeled with complete confidence. Transformative advances in disk imaging provide the first direct information on the initial conditions for planet formation, while exoplanet data has motivated new formation models that are faster, more efficient, and lead to a more diverse set of architectures than their Solar System inspired forebears. Much remains to be learned, and I close with a personal, incomplete list, of open problems.
@@ 1 Introduction
This review provides an introduction to this standard model, and serves as an overview of the following Chapters which discuss planet formation theory in detail. I introduce the physical processes relevant for planet formation, and summarize how they fit together in our consensus understanding of how planets form. Although a plausible outline of planet formation theory has been known for decades, existing models are surely not the last word: despite transformative advances in exoplanet detection and disk imaging there remain few direct observations of planet formation processes. It is conceivable that are fundamental gaps in our current theory. Even within the established paradigm, some phases are known to be poorly understood, and I conclude by listing some open questions that I hope to see resolved soon.
@@ 2 Disk formation and evolution
Although limited grain growth may occur during star formation prior to disk formation (Kwon et al. 2009; Marchand et al. 2023), the vast majority of planet formation processes occur within protoplanetary disks. Disks are flattened near-equilibrium fluid structures, whose properties are determined observationally to varying (mostly low) levels of precision. Typical stellar accretion rates scale roughly as $\dot{M}\sim 10^{-8}(M_{*}/M_{\odot})^{2}M_{\odot}{\rm yr}^{-1}$, with a measurement precision of $\sim 0.35\ {\rm dex}$ and a spread at fixed stellar mass $M_{*}$ of more than an order of magnitude (Manara et al. 2023). The disk lifetime is around 3 Myr (Haisch et al. 2001), though the absolute ages of Young Stellar Objects are notoriously hard to determine (Soderblom et al. 2014). The relative ages of young clusters are more reliable. Disk masses can be estimated starting from various observables, including the dust continuum flux (Ansdell et al. 2016; Tychoniec et al. 2020), molecular line emission in commonly observed sub-mm tracers (CO and its isotopologues, sometimes corrected using additional species such as N2H+; Trapman et al. 2022), HD emission (McClure et al. 2016), and (at the high mass end) rotation curves (Martire et al. 2024; Andrews et al. 2024). All of these methods have drawbacks – they either work only for a subset of disks or have possibly large uncertainties – such that a reliable empirical model for disk evolution remains elusive. Commonly adopted initial conditions for planet formation studies include a disk gas surface density profile that scales with radius as $\Sigma\propto r^{-1}$, an initial gas to dust ratio of $10^{2}$, and a mass within 100 au of $10^{-3}-0.1\ M_{*}$, but this is more a matter of art and convention than anything based on solid data. Despite its well-known shortcomings, the Minimum Mass Solar Nebula (Weidenschilling 1977b; Hayashi 1981) also lives on in some computers.
@@
where $\rho$ is the density, $\Omega_{\rm K}$ the Keplerian angular velocity, and we have assumed that $z\ll r$, the radial distance to the star. For a vertically isothermal disk with sound speed $c_{s}$, $P=\rho c_{s}^{2}$, and the gas density declines as a gaussian,
@@
Except in cases of unusually high accretion rates, or when close to the star, the heating of protoplanetary disks occurs primarily from the absorption and re-radiation of stellar luminosity (Kenyon and Hartmann 1987). Irradiation leads to a radial temperature profile approximately given by a power-law, $T\propto r^{-3/7}$ (Chiang and Goldreich 1997), and a disk that flares, such that $h/r$ increases with radius. Typically, $0.10.03<h/r<$ 0.1.
In the radial direction, the force balance involves gravity, centrifugal force, and a pressure gradient,
@@
The pressure gradient term is order $(h/r)^{2}$ smaller than the gravitational one, so to leading order $\Omega(r)=\Omega_{\rm K}$. The small correction – which usually but not always leads to sub-Keplerian rotation – is nonetheless extremely important. The dynamics of dust particles (Whipple 1972; Weidenschilling 1977a), the formation of planetesimals (Youdin and Goodman 2005), and the interpretation of observed disk sub-structures (Andrews 2020) are all thought to hinge on this $\sim$1deviation from strict Keplerian motion.
Figure 1: An illustration of the main processes that lead to evolution of the gaseous component of protoplanetary disks. Observational evidence and theoretical expectations suggest a typical scenario in which the initial properties of disks are determined by infall and self-gravity, accretion is largely driven by torques from magnetohydrodynamic (MHD) winds, and dispersal is a consequence of photoevaporative and MHD outflows. Depending on the star formation environment, however, it is possible for streamers to be dynamically important at later times, or for strong external photoevaporation or stellar flybys to truncate disks early on. Turbulence, even if it is a sub-dominant process for disk evolution, remains critically important for many phases of planet formation.
@@
  • Processes that redistribute angular momentum within the disk, while conserving disk mass (up to the fraction of gas that is physically accreted by the star). Lesur et al. (2023) review the long laundry list of possible processes, most of which are associated with linear instabilities of disk flow. Self-gravity (Toomre 1964; Kratter and Lodato 2016), the magnetorotational instability (Balbus and Hawley 1991) as (strongly) modified by non-ideal magnetohydrodynamic effects, and the Vertical Shear Instability (Nelson et al. 2013) are the leading suspects. These processes generate turbulence, which has distinctive (and potentially observable) characteristics on large scales.
  • Loss of mass and / or angular momentum through disk winds. Photoevaporation (Hollenbach et al. 1994) – in which the surface of the disk is heated by high-energy photons to the point where it escapes in a thermal wind – is guaranteed to occur at some level (Alexander et al. 2014; Ercolano and Pascucci 2017). It is a pure mass-loss process. Magnetized winds (Blandford and Payne 1982; Pudritz and Norman 1983; Bai and Stone 2013), on the other hand, remove gas that has more angular momentum than the disk at the launch point, simultaneously dispersing the disk while driving accretion (Suzuki and Inutsuka 2009; Armitage et al. 2013)
Figure 1 illustrates these possibilities. In addition to global evolution, many of these processes can also plausibly generate rings, spirals, and other sub-structures in disks.
The state-of-the-art in direct numerical simulation (Xu and Kunz 2021; Mauxion et al. 2024) allows disks to be followed for about $10^{5}\ {\rm yr}$, roughly the duration of the Class 0 phase of Young Stellar Object evolution. Longer duration evolution, and studies that seek to compare theory against observed populations of disks, necessarily require simpler models. The minimal goal of one-dimensional models is to predict $\Sigma(r,t)$, though it is also possible to evolve the disk central temperature (this is essential for modeling disk outbursts; Bell and Lin 1994; Armitage et al. 2001; Zhu et al. 2009). The vertical magnetic flux threading the disk, $B_{z}(r,t)$, is another important dynamical quantity that in principle should be evolved in models that include magnetized winds or the magnetorotational instability. The simplest case is a disk that evolves solely due to local turbulent redistribution of angular momentum. The appropriate one-dimensional description is a diffusion equation (Lynden-Bell and Pringle 1974; Pringle 1981),
@@
in which $\nu$ is an effective viscosity that is generated by turbulence. If $\nu$ is approximated by a time-independent power-law function of radius, this equation admits a similarity solution (taking the form of a time-varying exponentially truncated power-law for the surface density) that has been widely used for disk modeling (e.g. Hartmann et al. 1998). More generally, the viscosity can be written in terms of the local scale height $h$ and disk sound speed $c_{s}$ as (Shakura and Sunyaev 1973, the “$\alpha$-prescription”),
@@
with $\alpha$ being a dimensionless parameter that characterizes the efficiency of angular momentum transport.
@@
Wind-driven evolution can be modeled within an analogous framework, by adding sink and advective terms to the right hand side of equation (5) that represent the effects of mass and angular momentum loss on the underlying disk. The detailed form of the additional terms depends on the choice of wind model. Tabone et al. (2022), by introducing an “$\alpha$-like” parameter for the wind torque, derived an analytic generalization of the Lynden-Bell and Pringle (1974) solution for one choice that captures the generic effects of combined viscous / wind-driven disk evolution. Chambers (2019) derived a solution that includes the effects of irradiation and viscous heating on the evolving disk structure. These analytic models are useful for rapid population modeling. The effects of infall of gas onto the disk (Lin and Pringle 1990), and the long timescale evolution of the net magnetic flux (Lubow et al. 1994), also have one-dimensional descriptions, though the more complex physics of these processes is harder to capture reliably in a reduced model.
@@
Determining the relative importance of these evolution processes, from either theoretical or observational arguments, is a major challenge. It has often been assumed that the ordering from top to bottom in Figure 1 maps onto a temporal sequence, in which disks form from infall on a time scale of $\sim 10^{5}\ {\rm yr}$, evolve initially as massive self-gravitating structures and then more slowly under the action of viscosity and winds, and are finally dispersed by photoevaporation. This remains a plausible paradigm. Observations of molecular line broadening (Flaherty et al. 2017), and of the thickness of dust layers (Villenave et al. 2022, 2025), imply low levels of turbulence in many (though not all; Flaherty et al. 2024) of the best-studied disks. The inferred values of $\alpha$ in the “weakly turbulent” disks – typically $\alpha\sim 10^{-4}$ or lower – are broadly consistent with strongly damped MHD turbulence (described as “dead zone” models; Gammie 1996) and weak purely hydrodynamic driving mechanisms. As a consequence, magnetized disk winds rather than turbulence are considered to be dominant in most recent models. There remains room, however, for larger deviations from conventional wisdom, in particular models in which infall is dynamically important at late as well as early times. Indeed, simulations of star formation (reviewed by Kuffmeier 2024) have long found that encounters and infall drive a highly dynamic mode of disk evolution (Seifried et al. 2013; Bate 2018).
@@
A primary result of ALMA (Atacama Large Millimeter/submillimeter Array) observations has been the discovery that protoplanetary disks typically show pronounced substructure, in the form of departures from the simplest model of an axisymmetric disk with a monotonically decreasing profile of intensity (Andrews et al. 2018). Substructure can be detected most economically in dust continuum imaging, but is also seen in molecular line data that directly traces the gas. Rings (ALMA Partnership et al. 2015), crescent-shaped features (van der Marel et al. 2013), and spirals (Pérez et al. 2016) are all observed, roughly in decreasing order of frequency (for reviews, see Andrews 2020; Bae et al. 2023). Substructure is also detected in scattered light images of disks (Avenhaus et al. 2018), most prominently around more massive stars. Scattered light substructure arises from small grains in the disk atmosphere, and is often harder to interpret than sub-mm images that trace regions closer to the mid-plane. Resolution limitations mean that all of the observed substructure involves orbital radii and spatial scales $r>{\rm au}$; typically the impact on planet formation would be in the outer disk at $r>10$ au, well beyond the water snowline. Different physical processes may very well lead to the existence of cavities and rings at much smaller, sub-au scales, including from the interaction of a stellar magnetosphere with the disk (Zhu et al. 2024) and from the onset of stronger turbulence due to the magnetorotational instability at temperatures where the alkali metals are thermally ionized (Gammie 1996). These structures could be central to our understanding of close-in planet formation (Chatterjee and Tan 2014).
@@
The proximate cause of currently observed substructure in dust continuum images is understood to be the aerodynamic concentration of particles in local gas pressure maxima (or, in weaker cases, local pressure enhancements over a smooth background). This process – often described imprecisely as “trapping” – was foreseen by Whipple (1972) (look at his amazingly prescient Figure 1). The pressure maxima can be in form of axisymmetric rings (Haghighipour and Boss 2003), as in the original Whipple proposal, vortices (Barge and Sommeria 1995), or spiral waves (Rice et al. 2004), matching the main classes of observed structures. Ring-like dust structures are also predicted to form without the need for underlying gas perturbations in the vicinity of condensation / sublimation fronts (Zhang et al. 2015), but this effect does not appear to be responsible for most of the substructure that is currently being seen.
@@
While the immediate cause of dust substructure is reasonably clear, the same cannot be said for why gas disks develop pressure maxima in the first place. A popular hypothesis is that the source of substructure is gravitational perturbations from already-formed planets. Planets launch spiral density waves, and open gaps whose edges – which are approximately axisymmetric pressure maxima – can become unstable to vortex generation through the Rossby Wave Instability (Lovelace et al. 1999). Planet-disk interaction models generate disk substructure that is often a good match to what is actually observed (Zhang et al. 2018), and are surely responsible for some substructure. That said, simulations of disk turbulence and magnetized winds also commonly lead to the formation of rings (Johansen et al. 2009; Simon and Armitage 2014; Suriano et al. 2018; Riols and Lesur 2019) and vortices, while self-gravity of course forms spirals. Infall can also generate substructure (Kuznetsova et al. 2022). As a result, the interpretation of substructure is uncertain. For the purposes of this review, the key takeaway is that the existence of substructure has profound implications for planet formation irrespective of the interpretation. If ringed disks almost always host planets, that means that a substantial population of ice or gas giant planets is able to form surprisingly quickly at orbital separations that are as large as $\sim 50~{}{\rm au}$. If substructure is instead a byproduct of disk transport processes, then the observed concentration of solids that results modifies the evolution of all subsequent phases of planetary growth.